OpenStudy (monteskellie):

Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jennifer to complete this job on her own.

OpenStudy (mathmate):

Again, reduce everything to how much they can work in one hour. Let C=number of hours Carl takes to do the job J=number of hours Jennifer takes to do the job Then "Carl can paint a room 3 hours faster than Jennifer can." means 1/C-1/J = 1/3 ...........(finishes three hours faster) " If they work together, they can complete the job in 2 hours." means 1/C+1/J=1/2 This looks like a difficult equation to solve, but whenever you have the special case like a+b=S a-b=D Then a=(S+D)/2, and b=(S-D)/2. Applying the shortcut, 1/C=(1/2+1/3)/2=(5/6)/2=5/12 1/J=(1/2-1/3)/2=1/12 So Jennifer will take 1/(1/12)=12 hours to finish the job. Carl will take 1/(5/12)=2.4 hours to finish the job. check: 5/12-1/12=4/12=1/3 .... yes, each hour Carl will finish 1/3 more than Jenn. 5/12+1/12=6/12=1/2......yes, the two work together will finish in two hours.

OpenStudy (monteskellie):

I believe it is 6 but not sure

OpenStudy (mathmate):

Take the time to read through my solution and understand every step. If you have questions, feel free to ask. I or someone else will be pleased to help you!

OpenStudy (monteskellie):

I am still a little confused by I will try to figure it out

OpenStudy (welshfella):

you can also find it by solving 1 / (3 - x) + 1/x = 1/2 where x is the number of hours taken by jennifer

OpenStudy (welshfella):

sorry thats not correct it should be 1 / (x - 3) + 1/x = 1/2

OpenStudy (welshfella):

this will give you 2 values of x but only one is possible.